OpenStudy (anonymous):

How do I find three arithmetic means between -2 and 12?

OpenStudy (amistre64):

a1 a2 a3 a4 a5 -2 12 the 3 terms in the middle are the means. arithmetic mean would have us defined an arithmetic sequence of the format: an = a1 + d(n-1)

OpenStudy (amistre64):

as such a5 = a1 + d(4) a5-a1 = d(4) a5-a1 ----- = d ; would be the difference between terms 4

OpenStudy (amistre64):

or, a3 would be the normal use of the term average, between -2 and 12 12-2 = 10/2 = 5 a1 a2 a3 a4 a5 -2 5 12 and a2 is the average of a1 and a3; similar concept for a4